### v1

# Disentangled Representation with Cross Experts Covariance Loss for Multi-Domain Recommendation

Abstract.
Multi-domain learning (MDL) has emerged as a prominent research area aimed at enhancing the quality of personalized services. The key challenge in MDL lies in striking a balance between learning commonalities across domains while preserving the distinct characteristics of each domain. However, this gives rise to a challenging dilemma. On one hand, a model needs to leverage domain-specific modules, such as experts or embeddings, to preserve the uniqueness of each domain. On the other hand, due to the long-tailed distributions observed in real-world domains, some tail domains may lack sufficient samples to fully learn their corresponding modules. Unfortunately, existing approaches have not adequately addressed this dilemma. To address this issue, we propose a novel model called Crocodile, which stands for Cross-experts Covariance Loss for Disentangled Learning. Crocodile adopts a multi-embedding paradigm to facilitate model learning and employs a Covariance Loss on these embeddings to disentangle them. This disentanglement enables the model to capture diverse user interests across domains effectively. Additionally, we introduce a novel gating mechanism to further enhance the capabilities of Crocodile. Through empirical analysis, we demonstrate that our proposed method successfully resolves these two challenges and outperforms all state-of-the-art methods on publicly available datasets. We firmly believe that the analytical perspectives and design concept of disentanglement presented in our work can pave the way for future research in the field of MDL.
Multi-domain Learning; Disentangled Representation Learning; Dimensional collapse
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Crocodile: Cross Experts Covariance for Disentangled Learning in Multi-Domain Recommendation; October 21-25, 2024; Boise, Idaho, USAisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
((a)) Dillema of Interest Diversity v.s. Collapse Prevention
((b)) gAUC in Kuairand1k and AliCCP
Figure 1. The proposed Crocodile successfully resolves the dilemma of interest diversity v.s. collapse prevention and achieves the best performance regarding gAUC.
Online recommendation systems (Broder, 2008) are now essential tools for businesses to effectively reach their target audiences, servicing billions of internet users daily. Empowered by advanced algorithms and vast traffic, these systems play a crucial role in modeling user experiences and generating revenue for advertisers. However, traditional methods still face substantial challenges due to the dynamic nature of online platforms, which are marked by diverse user preferences across domains. To address these challenges, researchers have increasingly turned to multi-domain learning (MDL) (Chang et al., 2023; Zhou et al., 2023; Sheng et al., 2021; Yang et al., 2022; Li et al., 2023) as a promising approach to enhance the performance by involving information from multiple related domains.
Unfortunately, MDL for recommendation often confronts the challenge of preference conflict, where a user might exhibit divergent behaviors across domains. To remedy this, the MDL community aims to develop domain-specific components, including STAR (Sheng et al., 2021), HiNet (Zhou et al., 2023), and PEPNet (Chang et al., 2023) have progressively integrated domain-specific components.
Interestingly, we observe a trend in MDL methods towards the development of bottom-level domain-specific modules. In this work, we uncover the mechanisms behind this trend. To the best of our knowledge, this is the first study to reveal the mechanism of increasing interest diversity to resolve preference conflict. Furthermore, we propose a novel metric, the Diversity Index (DI), to quantify interest diversity, as shown in Fig 1(a).
Based on our discovery, we notice that domain-specific embedding (Su et al., 2023; Guo et al., 2023a) incurs the highest diversity and hence should be able to resolve preference conflict. However, naively deploying domain-specific embeddings does not turn out to be fruitful, we observe negative transfer in domains with fewer data. Delving further, we find out the core issue is sub-optimal embeddings for the domain with less data, i.e., they are dimensionally collapsed. [^1] We adopt the note of dimensional collapse in the work of Chi et al. (2022) and utilize information abundance (IA) (Guo et al., 2023b) to quantify dimensional collapse.
As shown in Fig. 1(a), our analysis demonstrates that existing works struggle to maintain these two properties: maintaining diversity and avoiding dimensional collapse, thus resulting in the sub-optimal performance shown in Fig. 1(b). As a result, it remains an open question whether we can incorporate both properties.
To this end, we propose a novel a novel Cross-experts Covariance Loss for Disentangled Learning model (Crocodile) framework, which we build this Covariance Loss (CovLoss) to alleviate embedding collapse while in the meantime enabling learning diverse user interests across various domains. To further unleash the power of Crocodile, we develop a novel Prior Informed element-wise Gating (PeG) mechanism to work with CovLoss.
Empirical analysis on two public datasets confirms that Crocodile achieves state-of-the-art results. Further analysis also demonstrates that Crocodile can simultaneously mitigate embedding collapse and learn diverse representations. Ablation experiments reveal that each proposed component contributes positively to the overall performance. Our contributions can be briefly summarized as follows:
  • We discover the mechanism of previous works in maintaining diversity, especially the trend of focusing more on bottom-level modules. We propose a novel metric, Diversity Index (DI), to quantify the diversity and conclude that the trend of previous works is optimizing to increase DI.
  • We reveal that existing works fall into the dilemma of maintaining diversity and avoiding collapse, hence resulting in inferior performance.
  • We propose the Crocodile approach, which successfully balances both properties. Empirical evidence on two public datasets demonstrates that Crocodile achieves state-of-the-art performance and resolves the aforementioned dilemma through comprehensive analysis.
2. Related Work
2.1. Multi-domain Recommendation
Multi-domain learning (MDL) is a promising topic that has garnered attention from industrial recommendations. Many MDL methods have made great efforts to model complex multi-domain interests. Among them, MoE-based methods are widely applied in industrial recommendation systems, such as Taobao (Sheng et al., 2021) , Meituan (Chang et al., 2023) and Kuaishou (Zhou et al., 2023).
STAR (Sheng et al., 2021) utilizes the dot product of domain-specific and domain-shared weight matrices to capture domain-specific interests. HiNet (Zhou et al., 2023) proposes two expert-like components to capture cross-domain correlations and within-domain information explicitly. PEPNet (Chang et al., 2023) introduces PPNet and EPNet, which incorporate domain information through gating. These methods often innovate from the perspective of making more bottom-level modules be domain-personalized within the model.
Recently, STEM (Su et al., 2023) proposed a multi-embedding (ME) structure in multi-task learning (MTL) and step forward in allocating more bottom-level domain-personalized structure, i.e., embedding. It achieved better results compared to both single and multi-embedding versions of MMoE (Ma et al., 2018b) and PLE (Tang et al., 2020) in MTL experiments. However, STEM fails in the multi-domain recommendation. We will discuss the reasons in detail in the subsequent sections.
Besides, some other paradigms are applied for multi-domain recommendation. AdaSparse (Yang et al., 2022) and ADL (Li et al., 2023) applied adaptive learning mechanism. EDDA (Ning et al., 2023a), COAST (Zhao et al., 2023) and UniCDR(Cao et al., 2023) provided graph learning-based solutions.
2.2. Disentangled Representation Learning
In the realm of disentangled representation learning (DRL), methods based on spatial or statistical relationships have been proposed.
On the one hand, methods based on spatial relationships achieve disentangling by manipulating the distances or angles between representations. OLE (Lezama et al., 2018), dot product(Wu et al., 2021) and cosine similarity(Liu et al., 2021) push representations towards orthogonal. GAN-based models (Chen et al., 2016; Yu and Welch, 2021; Khrulkov et al., 2021) promote the discrimination among representations. However, they are computationally expensive and not conducive to online recommendation.
On the other hand, statistical methods view representations as instances drawn from high-dimensional distributions and design loss based on statistical measures. Distance correlation (dCorr)(Liu et al., 2022a) encourages lower dCorr between different factors. ETL (Chen et al., 2023) models the joint distribution of user behaviors across domains with equivalent transformation. VAE-based methods (Ma et al., 2019; Wang et al., 2022; Liu et al., 2022b) encourage higher KL divergence between dimensions. However, they often assume a normal prior distribution, which is challenging to satisfy in the context of sparse interactions (Rendle, 2010), especially under sample volume imbalance of multi-domain recommendation.
Besides, some methods integrate both aspects mentioned above. VICReg (Bardes et al., 2021; Shwartz-Ziv et al., 2024) employs covariance and variance as loss components within representations to promote independence across dimensions and prevent collapse while utilizing Euclidean distance between representations to ensure semantic similarity among representations of different patches of an image.
2.3. Dimensional Collapse
Dimensional collapse(Chi et al., 2022) is defined as representations that end up spanning a lower-dimensional space, which is evident through the pronounced imbalance in the distribution of singular values. Chi et al. (2022) found that SMoE structure (Fedus et al., 2022; Lepikhin et al., 2020) may lead to dimensional collapse due to the influence of gating. The work of (Guo et al., 2023b) claimed that this problem is also observed in popular models (e.g., DCN V2 (Wang et al., 2021) and Final MLP (Mao et al., 2023)). To measure the collapse, they proposed the Information Abundance (IA) (Guo et al., 2023b) metric, defined as the ratio between the sum and the maximum l1-norm value of the target singular values. A small IA indicates severe insignificance of tail dimensions and potential dimensional collapse.
Beyond a single perspective, some methods (Hua et al., 2021) built the bridge between representation disentanglement and dimensional collapse prevention. It also claimed feature decorrelation helps alleviate dimensional collapse by reducing strong correlations between axes and promoting more diverse and informative representations.
3. On the Challenges of Multi-Domain Recommendation
As we have discussed in Sec. 1, existing works fall into the dilemma of avoiding dimensional collapse and maintaining interest diversity. In this section, we will discuss these two properties thoroughly.
3.1. Preventing Dimensional Collapse
According to Fig 1(a), domain-specific embeddings demonstrate its effectiveness in capturing diverse user interests. However, the naive adoption of domain-specific embeddings does not yield the expected outcomes. In fact, the methods with domain-specific embeddings, i.e., ME-PLE (Tang et al., 2020; Su et al., 2023) and STEM (Su et al., 2023), both performance deteriorate in the domain with less data. For example, in domain S6 of Kuairand1k (Gao et al., 2022) dataset, the ME-PLE and STEM show -0.4% and -0.1% lower AUC than the single domain method. This phenomenon has sparked our deep investigation into its underlying mechanism.
Intuitively, the domain-specific embeddings are trained solely on data from the corresponding domain, which results in insufficient samples for domains with less data. Insufficient training data fails to provide adequate variance for high-dimensional representations, resulting in high-dimensional representations residing within a lower-dimensional subspace, called dimensional collapse (Guo et al., 2023b; Chi et al., 2022; Hua et al., 2021).
To measure the collapse, we adopt the IA metric (Guo et al., 2023b), which is defined as $IA(E)=||\sigma||_{1}/||\sigma||_{\infty}$, where $E$ is the embedding of feature (e.g., item id) and $||\cdot||_{1},||\cdot||_{\infty}$ represent the $l1$-norm and $\infty$-norm, respectively.
Remark 1.
Intuitively, singular values $\sigma_{i}$ represent the significance of $i$-th singular value, and IA captures the balance of the singular value distribution. Thus, a small IA indicates severe insignificance of tail dimensions and potential dimension collapse.
((a)) $log(IA)$ of Experts
((b)) $log(IA)$ of Item Embeddings
Figure 2. Information Abundance ($log(IA)$) dynamics of STEM’s bottom layer of experts and item embeddings on Kuairand1k. High and Low denote the items with the highest and lowest frequencies, respectively.
Taking STEM as the example of methods applying domain-specific embedding, we analyzed its training IA dynamics of experts and embeddings in Fig. 2. In Kuairand1k, S1 and S6 are the domains with the most and least data. In Fig. 2(a), the S6-specific expert exhibits a lower IA compared to S1, indicating that domains with a smaller data volume are more susceptible to dimensional collapse at the expert level. In Fig. 2(b), we further categorized items into five groups based on their exposure frequency and observed the collapse of item ID embeddings for both the highest and lowest frequency groups. We discovered that domain S6 exhibits a lower IA than the overall S6, indicating a more severe collapse, regardless of the group. Within each domain, the low-exposure group consistently shows a lower IA than the high-exposure group. This further corroborates the notion that a lower exposure count (i.e., data volume) is more prone to dimensional collapse.
Both analytical and empirical evidence suggest that inadequate data of domain can lead to dimensional collapse of the domain-specific embedding, hence leading to inferior performance.
3.2. Interest Diversity
Intuitively, the ability to capture various user interests across domains can be directly characterized by the diversity of representations. When facing conflict preference across domains, a set of diverse representations is enabled to encode both prefer/non-prefer semantics and thus provide distinct vectors to towers of those conflict domains.
To this end, we propose a novel metric, the Diversity Index (DI), to quantify the representations’ ability to capture conflict preference across domains. Specifically, DI depicts the ratio of samples that contain both large and small norm of representations within the whole conflict samples, which is formulated as [^2] Here $AF$ and $SF$ are abbreviations for all forward and specific forward respectively, aligned with the concepts in  (Su et al., 2023).:
$\mathcal{D}_{con(i,j)}$ contains preference conflict samples between domain $i$ and $j$. $\mathbb{I}[\text{{condition}}]$ is the indication function, which is assigned to 1 if the condition is true; otherwise, 0. $M$ represents experts, $\mathbf{O}_{k}^{(p)}$ means the output of $p$-th expert and $k$-th sample, $\tau_{t}$ and $\tau_{b}$ are the thresholds of big and small l2-norm.
Remark 2.
As shown in Fig. 1(a), ME-MMoE (Guo et al., 2023a) exhibits significantly lower diversity compared to STEM (Su et al., 2023), and its improvement in diversity is relatively limited when compared to MMoE (Ma et al., 2018b). This suggests that the representations learned by ME-MMoE (Guo et al., 2023a) lack sufficient ability to solve preference conflict challenges.
Figure 3. Crocodile Architecture, which consists of a Multi-Embedding (ME) layer, a Cross-expert Covariance Loss (CovLoss), and a Prior informed Element-wise Gating (PEG) mechanism.
4. Methodology
We propose a multi-embedding architecture with Cross-experts Covariance Loss for Disentangled Learning (Crocodile) to resolve the dilemma, containing Covariance Loss (CovLoss) and Prior informed Element-wise Gating (PEG), as shown in Fig. 3.
4.1. Multi-Embedding Layer
The raw input features $x=\{x_{1},x_{2},\cdots,x_{F}\}$ contains $F$ fields. We employ a multi-embedding (ME) architecture (Su et al., 2023), which utilizes $M$ sets of embedding lookup tables. Each embedding looked up from each respective table is fed into the corresponding expert. Thus, in ME architecture, there are $M$ sets of experts. For a given $p$-th embedding table corresponding to $k$-th sample and $f$-th field, the looked-up embedding is represented as $e_{k,f}^{(p)}$. Thus, each $p$-th embeddings corresponded to $k$-th sample is:
(3) $e_{k}^{(p)}=[e_{k,0}^{(p)},e_{k,1}^{(p)},\cdots,e_{k,F}^{(p)}].$
4.2. Experts and Towers
Methods based on the Mixture of Experts (MoE) have been widely applied for multi-task (Ma et al., 2018b; Tang et al., 2020; Su et al., 2023) or multi-domain (Zhou et al., 2023; Chang et al., 2023) recommendation. Thus, we followed the same architecture and employed several sets of Multi-Layer Perceptrons (MLPs) as the experts and towers. Besides, we utilized a gating mechanism to compute the weights for the representation output from each expert. The output of the $p$-th expert can be represented as :
(4) $\mathbf{O}_{k}^{(p)}=MLP^{(p)}(e_{k}^{(p)}),$
where $\mathbf{O}_{k}^{(p)}\in\mathcal{R}^{1\times d}$ and ReLU serves as the inter-layer activation function. Through weighted by gating, the outputs $\mathbf{O}_{k}^{(\cdot)}$ of each expert are aggregated into a vector, denoted as $t_{k}^{(s)}\in\mathcal{R}^{1\times d}$, the input of $s$-th domain. The detail of gating will be introduced in Section 4.4.
Finally, each domain-specific tower outputs the final predicted value for the corresponding domain as :
(5) $\hat{y}_{k}^{(s)}=\sigma(MLP(t_{k}^{(s)})),$
where $\hat{y}_{k}^{(s)}$ is the predicted value of $k$-th sample in the $s$-th domain, $\sigma(\cdot)$ is the sigmoid function.
4.3. Cross-expert Covariance Loss
Many pioneering work (Shwartz-Ziv et al., 2024; Bardes et al., 2021; Ermolov et al., 2021) have highlighted the efficiency of tailoring the covariance into objective functions. However, they aimed to decorrelate the dimensions within each representation rather than disentangle among representations of all experts. To learn diverse user interests at the expert level, we propose the cross-expert covariance loss (CovLoss) to explicitly disentangle representations among experts, defined as:
(6) $\mathcal{L}_{Cov}=\frac{1}{d^{2}}\sum\limits_{p,q\in M\times M,p>q}||[\mathbf{O}^{(p)}-\overline{\mathbf{O}}^{(p)}]^{T}[\mathbf{O}^{(q)}-\overline{\mathbf{O}}^{(q)}]||_{1},$
where $\mathbf{O}^{(p)},\mathbf{O}^{(q)}\in\mathcal{R}^{N\times d}$ are the outputs of $p$-th and $q$-th experts, $\overline{\mathbf{O}}^{(p)}\in\mathcal{R}^{1\times d}$ is the average value of each dimension across all samples, and $||\cdot||_{1}$ is the $l1$-norm.
Finally, the overall loss function is defined as :
(7) $\mathcal{L}=\sum_{s=1}^{S}[\frac{1}{N_{s}}\sum_{k=1}^{N_{s}}\mathcal{L}_{BCE}(\hat{y}_{k}^{(s)},y_{k}^{(s)})]+\alpha\mathcal{L}_{Cov}$
where $N_{s}$ is the number of $s$-th domain’s samples, $\hat{y}_{k}^{(s)}$ is the label of $k$-th sample in domain $s$, and $\alpha$ is the weight of CovLoss.
4.4. Prior Informed Element-wise Gating Mechanism
CovLoss aims to disentangle different representations by minimizing the covariance between dimensions among different experts. However, the existing gating structure hinders the effectiveness of CovLoss. MMoE (Ma et al., 2018b) and PLE (Tang et al., 2020)’s gating mechanism assigns an identical weight to all dimensions within each expert; the dimensions within each expert are mutually constrained. PEPNet (Chang et al., 2023)’s proposed Gate Neaural Unit operates at the dimension level as well. However, it is designed to scale each dimension separately within the representation but not to disentangle these representations. Consequently, we propose an Element-wise Gating mechanism (EG), which implements weight control over corresponding dimensions between experts, as shown in Fig. 3:
(8) $t_{k}^{s}$ $=g^{s}_{k}\odot\mathbf{O}_{k},$
where $\mathbf{O}_{k}\in\mathcal{R}^{K\times d}$ is the concatnation across all experts respected to $k$-th sample, and $\odot$ is the element-wise multiplication.
Moreover, we are also concerned about the parameter burden of ME structure. In the original ME-MMoE structure, the number of embeddings is tied to the number of domains. This is because the gating is performed using $g^{s}_{k}=Softmax(e_{k}^{s}W_{g}^{s})$, which means the embedding $e_{k}^{s}$ is still related to a $s$-th gating. Thus, when it is deployed in situations with a large amount of domains, the scale of parameters becomes unmanageable. To address this issue, we proposed Prior Informed Gating (PG), which involves an independent set of domain-unrelated prior embeddings $r_{k}$ to control the gating. $r_{k}$ takes the prior features as input, denoted as $x^{prior}$. For a given $k$-th sample and $f$-th field, the looked-up prior embedding is $e_{k,f}^{prior}$. Thus, the input of each gating is:
(9) $r_{k}=[e_{k,0}^{prior},e_{k,1}^{prior},\cdots,e_{k,F^{\prime}}^{prior}],$
where prior embedding contains $F^{\prime}$ fields. we will select user ID, item ID, and domain ID as the prior embedding fields. We defined the PG mechanism as:
(10) $g^{s}_{k}$ $=Softmax(r_{k}W_{g}^{s}),$
where $r_{k}\in\mathcal{R}^{1\times l}$ is the prior embedding of $k$-th sample, $W_{g}^{s}\in\mathcal{R}^{l\times K\times d}$ and $Softmax(\cdot)$ is operated across the second dimension.
Furthermore, PG can be understood as an attention (Vaswani et al., 2017)-like mechanism, comprising query, key, and value components corresponding to Prior embedding $r_{k}$, gating weight matrix $W_{g}^{s}$ of $s$-th domain, and expert outputs $\mathbf{O}_{k}$, respectively. Its significance lies in weighting the learnt user interests through the interaction of user, item, and domain information provided by query and key.
4.5. Understanding the CovLoss
CovLoss decorates the representations of different experts in an element-wise fashion. Beyond statistic understanding, CovLoss also directly addresses the core objective of cross-domain modeling: enhance the diversity of representations.
For simplification, we consider the model involving only two experts, outputting $\mathbf{x}$ and $\mathbf{y}$ drawn from two high-dimensional distributions $\mathcal{X}$ and $\mathcal{Y}$, respectively, and the $i$-th element of $\mathbf{x}$ and $\mathbf{y}$ are $\mathbf{x}_{i}$ and $\mathbf{y}_{i}$, respectively. Following (Shwartz-Ziv et al., 2024), we assume the randomness comes from the input of the network.
To maximize the diversity of the outputs from both experts, the objective for each pair of samples can be expressed as the expectation of l2-norm difference:
(11) $\mathrm{E}[(||\mathbf{x}||_{2}$ $-||\mathbf{y}||_{2})^{2}]=\mathrm{E}[||\mathbf{x}||_{2}^{2}+||\mathbf{y}||_{2}^{2}-2||\mathbf{x}||_{2}||\mathbf{y}||_{2}]$
$=\sum_{i=1}^{d}[\mu_{xi}^{2}+\sigma_{xi}^{2}]+\sum_{j=1}^{d}[\mu_{yj}^{2}+\sigma_{yj}^{2}]-2\mathrm{E}[\sqrt{\sum_{j\geq i}(x_{i}y_{j})^{2}}]$
$\geq-2\mathrm{E}[\sqrt{\sum_{i=1}^{d}\sum_{j=1}^{d}(x_{i}y_{j})^{2}}]$
$\geq-2\mathrm{E}[\sum_{i=1}^{d}\sum_{j=1}^{d}\sqrt{(x_{i}y_{j})^{2}}]=-2\sum_{j\geq i}\mathrm{E}[x_{i}y_{j}]$
$=-2\sum_{i=1}^{d}\sum_{j=1}^{d}[\mu_{xi}\mu_{y_{j}}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}cov(x_{i},y_{j})}],$
where $x_{i},y_{j}\geq 0$ because the outputs from experts have been activated by ReLU before. According to the above Eq., minimizing $cov(x_{i},y_{j})$ can maximize the lower bound of diversity objective [^3] Although $\mu_{xi}\mu_{y_{j}}+cov(x_{i},y_{j})$ simultaneously incorporates both the first and second moments, they are nearly independent of each other in optimization.
5. Experiment
In this section, we provide comprehensive analyses to address the primary research questions (RQs) outlined below:
  • RQ1: How does the Crocodile perform compared with state-of-the-art recommenders in public datasets?
  • RQ2: Can Crocodile capture diverse user interests without introducing dimensional collapse issue?
  • RQ3: How does each proposed component contribute to performance gain?
  • RQ4: How sensitive is Crocodile to critical hyper-parameters and model capacity?
We employed two metrics to measure the model performance: AUC and gAUC. AUC measures the overall ranking performance, and gAUC evaluates the quality of intra-user item ranking.
5.1. Datasets
In this study, to examine the performance of the model in realistic recommendation, we carefully selected the Kuairand1k (Gao et al., 2022) and AliCCP (Ma et al., 2018a) datasets for our experiments. For Kuairand1k, we chose five domains with top-5 data volumes. For AliCCP, we selected all three domains. Tab. 1 is the statistics for the two datasets after low-frequency filtering. We removed features with fewer than 10 exposures and replaced them with default values.
Table 1. Statistics of Kuairand1k and AliCCP datasets.
#Sample #User #Item #Impr in Domain
Kuairand1k (Gao et al., 2022) 11,648,743 1,000 188,148 2.4M(S0) / 7.8M(S1) / 0.4M(S2) / 0.9M(S4) / 0.2M(S6)
AliCCP (Ma et al., 2018a) 85,316,519 241,080 499,321 32M(S1) / 52M(S2) / 0.6M(S3)
5.2. Baselines
To establish a performance benchmark for comparison, we will select comparative methods from both MTL and MDL for multi-domain recommendation. We apply the SharedBottom (Caruana, 1997) method in single/multi-domain contexts and subsequently introduce popular MTL methods such as MMoE (Ma et al., 2018b), PLE (Tang et al., 2020) (both Single-Embedding and Multi-Embedding (Su et al., 2023) versions), and STEM (Su et al., 2023). Besides, we choose representative MDL methods, including STAR (Sheng et al., 2021), PEPNet (Chang et al., 2023), HiNet (Zhou et al., 2023), and AdaSparse (Yang et al., 2022).
5.3. Performance Evaluation (RQ1 and RQ2)
Table 2. Performance on Kuairand1k dataset. Bold and underline highlight the best and second-best results, respectively. * indicates that the performance difference against the second-best result is statistically significant at p-value¡0.05.
Methods AUC gAUC
S0 S1 S2 S4 S6 Overall S0 S1 S2 S4 S6 Overall
Single Domain 0.73272 0.74202 0.80179 0.72505 0.75559 - 0.60336 0.59610 0.54581 0.55085 0.60562 -
MTL (SE) for MDR Shared Bottom 0.73535 0.74271 0.82033 0.72552 0.75945 0.78574 0.60335 0.59694 0.56337 0.56299 0.62500 0.66087
MMoE 0.73639 0.74279 0.82090 0.72793 0.76005 0.78594 0.60188 0.59658 0.56278 0.56325 0.62721 0.66057
PLE 0.73716 0.74271 0.81759 0.72737 0.75807 0.78565 0.60585 0.59710 0.56234 0.56276 0.62908 0.66093
MTL (ME) for MDR ME-MMoE 0.73842 0.74281 0.81897 0.72637 0.75985 0.78586 0.60649 0.59874 0.56620 0.56080 0.63001 0.66188
ME-PLE 0.73645 0.74240 0.80785 0.72727 0.75138 0.78499 0.61084 0.59717 0.55982 0.56170 0.61744 0.66075
STEM 0.73741 0.74267 0.80526 0.72776 0.75449 0.78527 0.60820 0.59701 0.56250 0.56365 0.62595 0.66096
MDL PEPNET 0.73732 0.74275 0.81396 0.72767 0.76239 0.78574 0.60812 0.59589 0.55435 0.55931 0.62733 0.66018
HiNet 0.73285 0.74264 0.81955 0.72698 0.76451 0.78563 0.60303 0.59830 0.56416 0.56267 0.62790 0.66131
STAR 0.71890 0.74071 0.81586 0.72203 0.75954 0.78336 0.58028 0.59396 0.55229 0.55896 0.62419 0.65652
AdaSparse 0.73445 0.74235 0.81575 0.71845 0.76151 0.78481 0.60653 0.59729 0.56316 0.56128 0.63150 0.66031
Crocodile 0.73874 0.74430* 0.81646 0.72838 0.75817 0.78683* 0.61075 0.60068* 0.56964 0.56366 0.63579 0.66373*
[…]
Table 3. Performance on AliCCP dataset. bold and underline highlight the best and second-best results, respectively. * indicates that the performance difference against the second-best result is statistically significant at p-value¡0.05.
Methods AUC gAUC
S1 S2 S3 Overall S1 S2 S3 Overall
Single Domain 0.61299 0.56544 0.61735 - 0.58077 0.53023 0.58739 -
MTL (SE) for MDR Shared Bottom 0.61925 0.59355 0.61952 0.61885 0.58908 0.57252 0.59076 0.58982
MMoE 0.61951 0.59324 0.61969 0.61935 0.58907 0.57269 0.59065 0.58992
PLE 0.62152 0.59476 0.61949 0.61931 0.58911 0.57256 0.59070 0.58947
MTL (ME) for MDR ME-MMoE 0.62271 0.59741 0.62179 0.62131 0.59002 0.57258 0.59148 0.59048
ME-PLE 0.62352 0.58727 0.62063 0.62132 0.58974 0.55777 0.59130 0.59049
STEM 0.61978 0.58187 0.61978 0.61907 0.58941 0.55101 0.59132 0.59036
MDL PEPNET 0.62273 0.59239 0.61988 0.62034 0.58851 0.56080 0.59085 0.58975
HiNet 0.62042 0.59515 0.62003 0.62022 0.58881 0.57191 0.59015 0.58970
STAR 0.62039 0.59204 0.61886 0.61949 0.58705 0.56527 0.58932 0.58855
AdaSparse 0.62192 0.59656 0.62173 0.62189 0.58819 0.57170 0.58972 0.58924
Crocodile 0.62448 0.59951 0.62192 0.62233* 0.59039 0.57085 0.59204* 0.59123*
We compared Crocodile with state-of-the-art recommenders in the Kuairand1k (Gao et al., 2022) and AliCCP (Ma et al., 2018a) datasets. All experiments are repeated 6 times and averaged results are reported, as shown in Tab. 2 and Tab. 3, respectively.
5.3.1. Overall Performance.
Our proposed Crocodile abrains the best overall AUC and gAUC in two datasets. In the Kuairand1k dataset, Crocodile outperforms the second-best method significantly by 0.09% and 0.19% in terms of overall AUC and gAUC, respectively. In the AliCCP dataset, Crocodile surpasses the second-best method significantly by AUC of 0.04% and gAUC of 0.07% than the second-best method.
Crocodile achieves larger AUC and gAUC almost for almost all domains in two datasets significantly compared to single domain method, with p-value¡0.05. Such positive transfer can be attributed to Crocodile’s better usage of multi-domain information and multi-embedding to enhance performance in each domain, without raising dimensional collapse.
Figure 4. Comparison of the singular value spectrum $log(\sigma)$ of item ID and user ID embeddings in Kuairand1k dataset. We reported $log(\sigma)$ of ME-PLE and STEM S6-specific embedding, while the original $log(\sigma)$ of single embedding or the average of other ME methods.
5.3.2. Comparison with MTL for MDR methods.
Multi-task learning (MTL) methods for Multi-Domain Recommendation (MDR) can be primarily categorized into single-embedding (SE) and multi-embedding (ME) paradigms (Su et al., 2023). We observed that some methods allocating domain-specific embedding underperform compared to SE methods, despite the fact they can capture diverse user interests in Fig. 5(a). For instance, in the comparison between PLE and ME-PLE on the Kuairand1k dataset in, Tab. 2, PLE achieved an overall AUC of $0.78565$ and gAUC of $0.66093$. In contrast, ME-PLE’s AUC and gAUC were $0.78499$ and $0.66075$, respectively, both lower than those of PLE.
To investigate the underlying reason, we conducted a domain-by-domain performance comparison and found that ME-PLE exhibited notably poorer results in the smallest domain, S6, with a negative transfer in AUC of -0.4% compared to the single domain. A similar negative transfer was also observed in STEM. We analyzed their embeddings via singular spectrum analysis (Guo et al., 2023b) and show the results in Fig. 4. We discovered that ME-PLE and STEM, in comparison to other methods, exhibited a clear dominance of top factors with much higher singular value over tail factors in the S6 embeddings, indicating the dimensional collapse issue.
In contrast, Crocodile’s embeddings exhibit a more balanced importance of singular values, thereby better supporting the high-dimensional representational space. Consequently, within domain S6, Crocodile outperforms ME-PLE by 0.68% in AUC and 1.8% in gAUC, and surpasses STEM by 0.37% in AUC and 0.98% in gAUC. Even though STEM learns the most diverse user interest, it still fails to surpass Crocodile due to dimensional collapse.
((a)) Crocodile vs MDL and MTL
((b)) Crocodile vs other Losses
Figure 5. Diversity Index (DI) of representations in the Kuairand1k dataset. There are five domains in Kuairand1k, resulting in 19 unique combinations. We plot the DIs on all combinations in descending order. We compared Crocodile with MDL and MTL methods in (a) and other disentanglement losses in (b).
5.3.3. Comparison with MDL methods.
The Multi-domain Learning (MDL) approaches typically evolve on a trend of progressively allocating domain-personalized components from top to bottom to capture more diverse user interest.
To empirically support this, we analyze the Diversity Index (DI) of both MDL and MTL methods in Fig. 5(a). Firstly, along with the trend from top to bottom, the DI of STAR (Sheng et al., 2021), HiNet (Zhou et al., 2023), and STEM (Su et al., 2023) reveals an increased trend. Secondly, methods with an ME structure typically exhibit higher DI than their SE versions, e.g., ME-PLE has higher DI than PLE.
However, as discussed in Section 5.3.2, the allocation of domain-specific embeddings can lead to the side effect of dimensional collapse, which in turn may result in a negative transfer. Consequently, we are curious as to whether methods that do not employ domain-specific embeddings can achieve favorable outcomes.
For example, HiNet (Zhou et al., 2023) only allocates domain-specific components at the expert level and displayed the best ability to prevent dimensional collapse in Fig. 4. It thus got the best AUC of 0.76451 in domain S6, in Kuairand1k. However, HiNet still falls in suboptimal performance in overall AUC and gAUC in both datasets compared to Crocodile. This is because of its low ability to capture diverse user interests in  5(a).
Conversely, Crocodile achieves the second-best DI, only slightly trailing behind STEM, which employs domain-specific embeddings but faces the issue of dimensional collapse. Consequently, Crocodile attains the best overall AUC and gAUC.
5.3.4. Comparison with Other Forms of Losses.
Through comparison with MTL and MDL approaches, we found that Crocodile captures the diverse interests of users (Fig. 5(b)) without introducing the issue of dimensional collapse (Fig. 4), thereby achieving the best overall AUC and gAUC scores in Kuairand1k and AliCCP.
Thus, can other disentangled losses also achieve favorable diversity and corresponding positive outcomes compared to those without them? We compare CovLoss with several other disentangling losses, including Dot Product (Dot) (Wu et al., 2021), Cosine Similarity (Cos) (Liu et al., 2021), Distance Correlation (dCorr) (Liu et al., 2022a) and OLE (Lezama et al., 2018) (F-norm form).
We first analyzed the ability to capture user diverse interests in Fig. 5(b). We observed that all disentangled losses could achieve higher DI than ME-MMoE-PeG, where ME-MMoE-PEG is the basic structure without any disentangled loss compared to others.
Table 4. Comparison of different disentanglement losses with the base model on overall AUC and gAUC in Kuairand1k. * indicates that the performance difference against base is significant at p-value¡0.05.
Losses AUC Lift(%) gAUC Lift(%)
Base 0.78611 - 0.66208 -
+Dot (Wu et al., 2021) 0.78667 0.06% 0.66285 0.08%
+Cos (Liu et al., 2021) 0.78634 0.02% 0.66263 0.06%
+dCorr (Liu et al., 2022a) 0.78621 0.01% 0.66214 0.01%
+OLE (Lezama et al., 2018) 0.78650 0.04% 0.66319 0.11%
+Importance (Shazeer et al., 2017) 0.78632 0.02% 0.66208 0.00%
+Trans5 (Chen et al., 2023) 0.78550 -0.06% 0.66017 -0.19%
+CovLoss
0.78683 0.07% 0.66373 0.17%
Among them, we found that dCorr (Liu et al., 2022a) loss achieves the smallest improvement of only 0.01% and 0.01% for AUC and gAUC lift, respectively in Tab. 4. This is coherent with its related poor diversity compared to other disentangled losses in Fig. 5(b). In contrast, ultilizing CovLoss achieves the best AUC and gAUC with the lift of 0.07% and 0.17% compared to the Base model. This evidence supports that the ability to learn diverse interests is important to improve performance.
Furthermore, we are interested in exploring whether the transfer of other types of loss functions, e.g., load balance (Shazeer et al., 2017) and transfer learning (Chen et al., 2023), could yield performance enhancements. As shown in Tab. 4, the Importance loss (Shazeer et al., 2017) resulted in a limited improvement, whereas Trans5, one of the effective versions of equivalent transformation losses in (Chen et al., 2023), had a detrimental effect. This can be attributed to that these loss functions do not explicitly promote diversity in learning. Instead, they focus on balancing the contributions of various experts or enhancing the robustness of representations, which does not meet the expected outcomes.
5.4. Ablation Study (RQ3)
Table 5. Ablation study in Kuairand1k and AliCCP. ME and SE are shorts for multi- and single-embedding, respectively. B and B+C are shorts for BCE Loss and BCE Loss with CovLoss, respectively. * indicates that the performance difference against Base is significant at p-value¡0.05.
Methods Components Kuairand1k Performance AliCCP Performance
Embedding Gating Loss AUC Lift (%) gAUC Lift (%) AUC Lift (%) gAUC Lift (%)
Base (MMoE) SE - B 0.78594 - 0.66057 - 0.61935 - 0.58992 -
Ablation Method ME - B 0.78586 -0.01% 0.66188 0.13% 0.62131 0.20% 0.59048 0.06%
ME PG B 0.78585 -0.01% 0.66208 0.15% 0.62170 0.24% 0.59080 0.09%
ME PEG B 0.78611 0.02% 0.66208 0.15% 0.62227 0.29% 0.59117 0.12%
ME PG B+C 0.78642 0.05% 0.66333 0.28% 0.62214 0.28% 0.59124 0.13%
Crocodile ME PEG B+C 0.78683* 0.09% 0.66373* 0.32% 0.62237* 0.30% 0.59134* 0.14%
In this section, we empirically test the effectiveness of our proposed components with an ablation study. We eliminated CovLoss, PeG, and ME one by one and compared their performance on Kuairand1k and AliCCP, as shown in Tab 5.
All three components are critical since removing either of them leads to a significant performance drop on both datasets. In particular, CovLoss is more critical in the Kuairand1k dataset, while CovLoss and EG are equally important in AliCCP.
Table 6. gAUC metric along with different numbers of embeddings (#Emb) on ME-MMoE-PG and Crocodile. * indicates that the performance difference against ME-MMoE-PG is significant at p-value¡0.05.
#Emb 2 3 4 5 6 7 8 9 10
ME-MMoE-PG 0.66138 0.66157 0.66206 0.66232 0.66203 0.66196 0.66286 0.66279 0.66285
Crocodile 0.66198* 0.66284* 0.66324* 0.66373* 0.6637* 0.66394* 0.66395* 0.66432* 0.66397*
Lift (%) 0.06% 0.13% 0.12% 0.14% 0.17% 0.20% 0.11% 0.15% 0.11%
5.5. Hyper-Parameters and Embedding Sets (RQ4)
In this section, we investigate the sensitivity of the model’s performance to hyperparameters and the number of embeddings. For hyperparameters, the primary parameter is the weight of CovLoss in Eq. 7, i.e., $\alpha$. Taking Kuairand1k as an example, we find that as long as $\alpha$ is not too small (below $2\times 10^{-5}$), Crocodile’s AUC and gAUC remain ¿0.78590 and ¿0.66174, thus maintaining its status as the second-best method.
Additionally, we analyzed the impact of model capacity on performance by adjusting the number of embedding sets. At the minimum number of MEs (i.e., 2-sets), Crocodile achieves a gAUC of 0.66198 and an AUC of 0.78595, which still surpasses other MTL and MDL methods, even when the MTL(ME) methods apply for 5-sets of embeddings. Moreover, as the number of embeddings increases, the performance gains of Crocodile over ME-MMoE-PG are robust ($+0.06\%\sim+0.20\%$), as illustrated in Tab. 6.
5.6. Computational Complexity
The Multi-Embedding (ME) paradigm has already been widely adopted in industry (Zhaok et al., 2021; Ning et al., 2023b; Su et al., 2023; Guo et al., 2023b; Pan et al., 2024). Compared to these ME architectures, our proposed Crocodile introduces the CovLoss upon it, which only influences the training process while not the inference at all. Its complexity is not only small and negligible when compared with the ME-based backbone, but it can be further reduced by sampling.
The complexity of CovLoss is $\frac{M(M-1)}{2}\times(d^{2}N)$, shown lower complexity than dCorr (Liu et al., 2022a), which is $\frac{M(M-1)}{2}\times(d^{3}N+d^{2}N)$, where $M$ and $N$ are the number of experts and samples, and $d$ is the dimension of expert output. Besides, the complexity of experts in MMoE structure is $M\times\sum_{i=1}^{l-1}Nd_{i}d_{i-1}$, where $d_{i}$ is the hidden layer dimension for i-th layer in all $l$ layers. This is much higher than CovLoss in practice.
Moreover, we can reduce the complexity of calculating covariance by sampling. We discovered that selecting as low as 32 samples yields no significant difference in computed covariance norms (p-value¿ 0.1) in the testing dataset, contributing as much as 99.2% complexity reducing, as we originally selected $N=1024$.
6. Conclusion
Multi-domain learning is vital in enhancing personalized recommendations. Existing MTL methods fail to learn both dimensional-robust representations and diverse user interests. We propose Crocodile, which employs a Covariance Loss (CovLoss) and Prior informed Element-wise Gating (PEG) based on the multi-embedding (ME) paradigm. Comprehensive experimental evaluations validate the effectiveness of Crocodile.